A firm purchases $4,562,500 in goods over a 1-year period from its sole supplier. The supplier offers trade credit under the following terms: 2/15, net 50 days. Davis finally chooses to pay on time (pay in the 50th day) but not to take the discount. We assume 365 days per year. What is the average level of the company’s free trade credit?

$187,500

Based on the information from Question 31, what is the effective annual cost of the firm’s costly trade credit?

17.6%

21.3%

23.5%

25.2%
Answer #1

50. Average Level of companys free trade credit = Total credit purchases in a year / 365 * Number of days credit availed

Average Level of companys free trade credit = 4,562,500 / 365 * 50 = 625,000

51. The company is not availing 2/15 net 50 discount

ie it has the option to take a discount of 2% and pay within 15 days ot pay the full amount in 50 days

Let us assume we have to pay 100 rs

We have 2 Options

Pay 98 on Day - 15

Or Pay 100 on Day 50

in a way by not paying on day 15 we have borrowed 98 for 35 days

Effective Annual cost of this borrowing = (100 / 98) ^ (365 / 35) - 1 = 23.50%